OpenStudy (anonymous):

If the expression x^2 + bx - 15 has a factor of (x - 3), what is the value of b?

TheSmartOne (thesmartone):

So (x-3)(x+_)=x^2+bx-15 So -3 * what=-15

TheSmartOne (thesmartone):

And then to find b you just add -3 and the number you found out that multiplies with -3 and equals to -15